OpenStudy (anonymous):

simplify the rational expression. 3x-6 over x-2

OpenStudy (anonymous):

\[\frac{ 3x-6 }{ x-2 } = \frac{ 3(x-2) }{ x-2 }\] the x-2 cancels out s what do you get as the answer??

OpenStudy (anonymous):

so u just get 3?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

do you understand the 1st step??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yw...(: